# A Cross-Scene Benchmark for Open-World Drone Active Tracking

### v1

@@ Abstract
Drone Visual Active Tracking aims to autonomously follow a target object by controlling the motion system based on visual observations, providing a more practical solution for effective tracking in dynamic environments. However, accurate Drone Visual Active Tracking using reinforcement learning remains challenging due to the absence of a unified benchmark, the complexity of open-world environments with frequent interference, and the diverse motion behavior of dynamic targets. To address these issues, we propose a unified cross-scene cross-domain benchmark for open-world drone active tracking called DAT. The DAT benchmark provides 24 visually complex environments to assess the algorithms’ cross-scene and cross-domain generalization abilities, and high-fidelity modeling of realistic robot dynamics. Additionally, we propose a reinforcement learning-based drone tracking method called R-VAT, which aims to improve the performance of drone tracking targets in complex scenarios. Specifically, inspired by curriculum learning, we introduce a Curriculum-Based Training strategy that progressively enhances the agent tracking performance in vast environments with complex interference. We design a goal-centered reward function to provide precise feedback to the drone agent, preventing targets farther from the center of view from receiving higher rewards than closer ones. This allows the drone to adapt to the diverse motion behavior of open-world targets. Experiments demonstrate that the R-VAT has about 400% improvement over the SOTA method in terms of the cumulative reward metric.
Figure 1: Example scenes of DAT benchmark. We construct 6 environments with 4 weather conditions for cross-scene and cross-domain generalization tasks, providing a unified benchmark for the VAT.
@@ 1 Introduction
Visual Active Tracking (VAT) aims to autonomously follow a target object by controlling the motion system of the tracker based on visual observations [73, 68]. It is widely used in real-world applications such as drone target tracking and security surveillance [20, 66, 70, 48]. Unlike visual passive tracking [4, 67, 28, 6, 9, 75, 60, 51], which involves proposing a 2D bounding box for the target on a frame-by-frame with a fixed camera pose, VAT actively adjusts the camera position to maintain the target within the field of view. Passive visual tracking often falls short in real-world scenarios due to the highly dynamic nature of most targets. Thus, VAT offers a more practical yet challenging solution for effective tracking in dynamic environments.
AD-VAT+ [73] D-VAT [18] AOT [33] DAT
@@
Dynamics
Randomness
@@
Recently, VAT methods have evolved into two main categories: pipeline VAT methods [34, 40, 14] and reinforcement learning-based VAT methods [18, 33, 17, 73]. Pipeline VAT methods employ a sequential framework where the visual tracking [29, 5, 56, 28] and control models are connected in series. Here, the object tracking model first processes the input image to estimate the target position and then the control model to generate the necessary control signals. While this modular design allows for clear task separation, these methods often require significant manual effort to label the training data, and the combination of modules requires extra effort for tuning and implementation. To address these issues, reinforcement learning-based VAT methods integrate visual tracking and control within a unified framework. These methods eliminate the need for separate tuning of the visual tracking and control modules by using a unified framework to map raw visual inputs directly to control actions. Therefore, the reinforcement learning-based VAT methods simplify system design and increase the efficiency of learning adaptive tracking behaviors in dynamic environments.
Unfortunately, achieving accurate drone visual active tracking with reinforcement learning remains challenging, partly for the following reasons. 1) Missing unified benchmark. Existing benchmark scenes are limited in scope, low in complexity, and few in number, and unable to adequately validate the performance of agents (see Tab. 1). Previous methods provide limited tracking targets and sensors, insufficient to establish benchmarks for different tasks. 2) Vast environments with complex interference. Open-world tracking involves large, dynamic environments with frequent interference. It brings significant challenges for agents to accurately follow targets. Training directly in these conditions often leads to slow convergence or difficulty in building robust tracking behaviors. 3) Complex targets with diverse motion behaviors. Open-world targets often exhibit complex, unpredictable behaviors, requiring agents to adapt to varied movements and orientations. Existing methods assume a fixed forward-facing view, leading to reward functions that inaccurately reflect tracking performance across different behaviors and perspectives.
Figure 2: A pipeline for drone visual active tracking.
To address the above limitations, we first propose a unified cross-scene benchmark for open-world drone active tracking (called DAT) that simulates the diversity and complexity of the real world as possible (see Fig. 1). Specifically, the DAT benchmark provides 24 visually complex scenes to validate the algorithms’ cross-scene and cross-domain generalization abilities. It offers comprehensive support for diverse tracking scenarios with 2 tracker types, 24 target types, and 6 different sensor types, with plug-and-play interfaces that facilitate the integration of custom robot models and controllers. To better replicate real-world conditions, DAT employs the webots simulation software [31] for high-fidelity modeling of realistic robot dynamics. It incorporates the Simulation of Urban Mobility (SUMO) [30] for managing target behavior, enabling diverse and efficient path modeling that surpasses traditional rule-based approaches. Second, we propose a novel drone visual active tracking with reinforcement learning method (called R-VAT), aiming to improve the performance of drone tracking targets in complex scenarios. Specifically, inspired by curriculum learning [58, 47, 35, 76], we propose a Curriculum-Based Training strategy that progressively improves agent performance in vast environments with complex interference. In addition, we design a goal-centered reward function to provide accurate feedback to the agent, enabling it to adapt to the target’s diverse motion behaviors. Unlike existing methods, this function is designed at the image level to prevent targets farther from the center of view from receiving higher rewards than closer ones.
Figure 3: Statistics and simulator component examples of the DAT. (a) Statistics on the 7 complexity aspects in DAT scenes. (b) Examples of challenging situations in various scenes. (c) Diagrams of the provided sensors. (d) Examples of available tracking targets.
To summarize, our contributions are as follows:
  • A comprehensive drone active tracking benchmark. We present the DAT benchmark, simulating 24 visually diverse, real-world environments with various trackers and targets, enabling rigorous cross-domain and cross-scene validation of algorithms.
  • A novel drone active tracking method. We propose the R-VAT, which leverages a Curriculum-Based Training strategy for tracking and a goal-centered reward function to enhance drone tracking performance in complex and dynamic environments.
  • Extensive experimental validation. Experiments validate the DAT benchmark usability, with R-VAT achieving about 400% improvement over the SOTA method in terms of the cumulative reward metric.
@@ 2 Task Definition of Drone Active Tracking
The DAT task seeks to train a model to control a drone for active target tracking in dynamic environments (see Fig. 2). Using visual and motion sensor data, the model learns actions to keep the target centered in view, ensuring robust performance across diverse scenarios.
Observation spaces. The target is initially positioned at the center of the camera’s field of view, and the observation space comprises data acquired from sensors (e.g., RGB images with a resolution of $84\times 84$).
Action spaces. There are two options for the action space: discrete action space and continuous action space. The discrete action space is a 7-dimensional vector, representing movements such as forward, backward, left translation, right translation, counterclockwise rotation, clockwise rotation, and stopping. The speed corresponding to each action must be set by the user. The continuous action space is a 4-dimensional vector that defines speed values in the forward, lateral, vertical, and yaw directions.
Success criterion of DAT task. We define a success criterion when the model can keep the target object, which is initially located at the center of the camera’s field of view, in the middle of the image for a long duration.
Challenges. Open-world drone active tracking presents substantial challenges primarily due to the scarcity of real-world data and the high costs and risks associated with real-world trial-and-error. This highlights the critical need for building complex and diverse simulation environments. Additionally, open-world scenes are characterized by high diversity and dynamic elements, which introduce complex interferences, further challenging the model’s robustness. Moreover, real-world tracking targets exhibit diverse and unpredictable behaviors, making it difficult for models to adapt. Therefore, improving the adaptability of models to track complex targets effectively is another key challenge.
@@ 3 DAT Benchmark with Diverse Settings
We develop the DAT benchmark (see Fig. 3) to evaluate the model’s cross-scene and cross-domain adaptability in 6 large outdoor environments with 4 weather conditions.
@@ 3.1 Diverse Scene Construction
The construction of the DAT scene aims to simulate the diversity and complexity of the real world. To enhance diversity, we select 6 categories of outdoor scenes and 4 weather conditions, covering almost all natural conditions that applications may face. To simulate the real-world complexity, we model 7 aspects: scene area, building density, color richness, road density, terrain density, tree density, and tunnel density. Specifically, the scene area refers to the extent of the scene, building density is the ratio of the number of buildings to the scene area, and color richness is the number of dominant colors in the scene. These three aspects primarily depict the complexity of the visual background. Road density is measured by evaluating the density of complex elements such as intersections, terrain density is determined by assessing the density of special terrain features like mountains in the scene. These two aspects primarily depict the complexity of tracking target behavior. The tree density and tunnel density are calculated as the ratio of the number of trees and tunnels to the scene area, and are used to measure the level of visual occlusion within the scene.
As shown in Fig. 3(a), the six different scenarios exhibit unique and realistic complexity across the seven aspects:
@@
•  Village scene spans 1.4 square kilometers. This scene features a mountainous terrain density of 20.1 and a tunnel density of 6.3, requiring the agent to predict the target’s movement when it is fully obscured by tunnels.
•  Downtown scene covers 0.8 square kilometers. The scene includes complex road elements and a very high building density of 304.9, challenging the agent’s tracking accuracy and obstacle avoidance capabilities.
•  Lake scene encompasses 1.6 square kilometers. The density of road elements is 68.1, and the richness of background colors is rated at 5, challenging the agent’s robustness across varying features and colors.
•  Farmland scene covers an area of 0.7 square kilometers. The color richness is rated at 5, with multiple color patches spread throughout the scene, posing a significant challenge to the agent’s adaptability to multi-color environments.
•  Desert scene covers 1.1 square kilometers. It includes a mountainous terrain density of 37.1 and a road element density of 31.0. Some roads are covered by sand, testing the agent’s adaptability to such challenging conditions.
The four weather conditions mainly test the agent’s cross-domain adaptability. Foggy reduces visibility, with a visibility range of 400m to 2000m. Night reduces brightness and light uniformity, and snow changes the color tone of the environment. The above 24 scenes can fully measure the performance of the agent active adaptation (see Fig. 1).
@@ 3.2 Various Trackers and Targets Construction
Drone Active Tracking in real-world applications involves various trackers and targets depending on the specific task. For example, drones are typically used as trackers, with automobiles as targets in security monitoring. The DAT benchmark provides diverse trackers and targets, making it adaptable for different use cases.
Tracker. The DAT benchmark supports two tracker types: drones and ground robots. The drone used is the DJI Matrice 100 [12], with a default flight controller [53]. The DAT benchmark provides visual and motion sensors (see Fig. 3(c)) that can obtain the state parameters of the drone relative to the world coordinate system. The drone position and velocity are obtained via GPS, angular velocity via gyroscope, Euler angles are obtained from the IMU and can be converted into quaternion, and acceleration via accelerometer. The LiDAR provides point cloud data. To better replicate real-world conditions, drones are equipped with a 3-axis gimbal, similar to those on commercial drones like the DJI Mini 3 Pro [13], enabling precise camera adjustments and adding complexity beyond static point trackers. For ground robots, users can select from various robot models. The benchmark employs webots simulation software to provide a high-fidelity dynamic engine, accurately modeling tracker movements, collisions, and interactions, crucial for evaluating tracking algorithms in challenging scenarios. Additionally, we provide common parameters for further task customization. See Appendix B for details.
Targets. The DAT benchmark includes five categories of targets: automobile, motorbike, pedestrian, wheel robot, and legged robot, providing a total of 24 tracking targets, each with built-in controllers (see Fig. 3(d)). To support custom robot designs, a plug-and-play interface is available, allowing users to easily integrate robot models and controllers into the benchmark. See Appendix B for details.
Target Management. Realistic target behavior and path diversity are essential for simulating authentic environments. The DAT benchmark uses the SUMO traffic simulator to manage all scene targets. For example, SUMO generates random vehicle paths and dynamically controls actions like acceleration, lane changes, and stops. It can also refresh vehicles based on configurations and randomize their types and colors, creating a high-fidelity traffic system.
@@ 4 VAT with Reinforcement Learning
In this paper, we primarily focus on visual active tracking (VAT), a core task within the DAT benchmark. We propose a drone visual active tracking with reinforcement learning method (called R-VAT), aiming to improve the performance of tracking targets in complex scenes. As shown in Fig. 2, we model drone active tracking as a Markov Decision Process (MDP) and train a Drone Agent to track a target in the open scene. At time step $t$, the Drone Agent takes the current state $s_{t}$ as input and selects an action. Then the reward is calculated via Eq. 2. The trajectories are collected to train the agent via the Curriculum-Based Training strategy.
@@ 4.1 MDP for Drone Active Tracking
Action $\boldsymbol{\mathcal{A}}$ is a set of discrete actions that the drone can take, including forward, backward, left translation, right translation, counterclockwise rotation, clockwise rotation, and stopping, respectively. At each time step, the Drone Agent selects an action $a_{t}\in\boldsymbol{\mathcal{A}}$ based on the state $s_{t}$ and actively controls the camera movement.
@@
Reward $\mathcal{R}(s_{t},a_{t})$ is the reward function. The reward we designed is constructed only at the image level. The details are given in Sec. 4.3.
Network structure of Drone Agent. Since the VAT task requires the agent to maintain long-term tracking, a dynamics model is essential for capturing temporal information. Therefore, similar to previous works [33, 73, 17], we select a CNN combined with a gated recurrent unit [11] network architecture as the backbone (see Appendix C).
4.2 Curriculum Learning for Agent Training
The DAT environment contains numerous dynamic targets and diverse obstacles, making it challenging for the agent to discover sufficient successful trajectories. This difficulty results in low convergence rates and limited performance. Progressively training the agent from simpler to more complex environments increases the percentage of successful trajectories, enhancing performance and accelerating learning for the final task [57]. Therefore, we propose a Curriculum-Based Training (CBT) strategy for tracking to optimize agent training in complex environments.
Training Algorithm. To address the challenges of learning in complex environments, we employ the Proximal Policy Optimization (PPO) [49] algorithm, which is widely recognized for its robustness and efficiency in complex continuous control tasks. Given the complexity of the environment, we adopt a multi-stage training strategy, which divides the model training into two stages to build the agent’s tracking capabilities progressively. The first stage $\mathcal{E}_{1}$, consists of a simplified environment where the target trajectory is a straight line, and there are no obstacles or visual interferences. The agent learns to keep the target centered in the image through the reward signal from Eq. 2. Once the reward obtained for agent in first stage $\mathcal{E}_{1}$ converges as:
$\small\frac{1}{T}\sum_{t=1}^{T}r_{t}\geq\eta,$ (1)
where $r_{t}$ represents the reward obtained by the agent at time step $t$, $T$ is the total number of time steps over which the reward is averaged, and $\eta$ is a predefined threshold indicating satisfactory performance. Once this condition is satisfied, the agent transitions to the second stage $\mathcal{E}_{2}$, where the agent encounters more varied tracking target movements and more complex visual information, such as tree occlusions, crosswalks, and poles. The goal for the agent is to develop stronger task generalization abilities based on task understanding gained in the first stage.
Data Augmentation. While simpler settings facilitate the agent’s learning of task objectives, they also heighten the risk of the agent rapidly converging to a locally optimal action distribution, which can undermine the exploration process. Consequently, implementing data augmentation is essential for enhancing the agent’s exploratory capabilities. This is achieved through the randomization of the drone’s initial position and orientation relative to the target, which necessitates a broader range of actions from the agent to maximize rewards. Moreover, to enhance the agent’s spatial perception ability, randomization is also introduced in its gimbal pitch angle. See Appendix C for details.
Figure 4: Schematic diagram of reward acquisition.

### v2

# Open-World Drone Active Tracking with Goal-Centered Rewards

@@ Abstract
Drone Visual Active Tracking aims to autonomously follow a target object by controlling the motion system based on visual observations, providing a more practical solution for effective tracking in dynamic environments. However, accurate Drone Visual Active Tracking using reinforcement learning remains challenging due to the absence of a unified benchmark and the complexity of open-world environments with frequent interference. To address these issues, we pioneer a systematic solution. First, we propose DAT, the first open-world drone active air-to-ground tracking benchmark. It encompasses 24 city-scale scenes, featuring targets with human-like behaviors and high-fidelity dynamics simulation. DAT also provides a digital twin tool for unlimited scene generation. Additionally, we propose a novel reinforcement learning method called GC-VAT, which aims to improve the performance of drone tracking targets in complex scenarios. Specifically, we design a Goal-Centered Reward to provide precise feedback across viewpoints to the agent, enabling it to expand perception and movement range through unrestricted perspectives. Inspired by curriculum learning, we introduce a Curriculum-Based Training strategy that progressively enhances the tracking performance in complex environments. Besides, experiments on simulator and real-world images demonstrate the superior performance of GC-VAT, achieving a Tracking Success Rate of approximately 72% on the simulator. The benchmark and code are available at https://github.com/SHWplus/DAT_Benchmark.
@@ 1 Introduction
Visual Active Tracking (VAT) aims to autonomously follow a target object by controlling the motion system of the tracker based on visual observations advat+ ; yuan2023active . It is widely used in real-world applications such as drone target tracking and security surveillance emran2018review ; XING2022102972 ; ZHANG2023119243 ; Schedl2021AnAD . Unlike passive visual tracking babenko2009visual ; ostrack ; SiamRPN0 ; SORT ; Chen_2023_CVPR ; 9573394 ; wen2021detection ; smeulders2013visual , which involves proposing a 2D bounding box for the target on a frame-by-frame with a fixed camera pose, VAT actively adjusts the camera position to maintain the target within the field of view. Passive visual tracking often falls short in real-world scenarios due to the highly dynamic nature of most targets. Thus, VAT offers a more practical yet challenging solution for effective tracking in dynamic environments.
Figure 1: A pipeline for drone VAT.
AD-VAT+ advat+ D-VAT dvat AOT e2e DAT
@@
Dynamics Simplified Full Physics
Target Behavior Policy-based Rule-based Rule-based Human-like
Scene Building Manual Manual Manual Digital Twin
@@
Recently, VAT methods have evolved into two main categories: pipeline VAT methods followanything ; fasttracker2 ; das2018stable and reinforcement learning-based VAT methods dvat ; e2e ; cvat ; advat+ . Pipeline VAT methods employ a sequential framework where the visual tracking siamrpn ; SiamFC ; SiamMask ; SiamRPN0 and control models are connected in series. The tracking model estimates the target position in the input image and the control model generates control signals. While this modular design allows for clear task separation, it often requires significant manual effort to label the training data, and the control module requires additional tuning for different scenes. To address these issues, reinforcement learning-based VAT methods integrate visual tracking and control within a unified framework. These methods eliminate the need for separate tuning of the tracking and control modules by using a unified framework to map raw visual inputs directly to control actions. Therefore, the reinforcement learning-based VAT methods simplify system design and increase the efficiency of learning adaptive tracking behaviors in dynamic environments.
Unfortunately, achieving accurate drone VAT with reinforcement learning remains challenging, partly for the following reasons. 1) Missing unified benchmark. Existing benchmark scenes are low in complexity, neglect tracker dynamics or rely on overly simplified models, making them inadequate to validate the agent performance (see Table 1). Previous methods e2e ; dvat ; airsim use rule-based target management, far from producing human-like target behaviors. Additionally, current 3D scenes are all manually constructed, leading to a heavy workload and limited scene number. 2) Vast environments with complex interference. Open-world tracking involves large, dynamic environments with frequent interference. In previous methods e2e ; dvat , trackers can only capture images from a fixed horizontal viewpoint. However, the fixed forward viewpoint captures excessive sky, reducing target-related visual information, especially for air-to-ground tracking tasks. Besides, since VAT goal is to keep the target at the image center, such viewpoint restricts the tracker to the same height as the target, severely limiting the perception and movement range. Moreover, training directly in complex conditions leads to slow convergence or difficulty in building strong behaviors.
To address the above limitations, we first propose DAT, the first open-world active air-to-ground tracking benchmark that simulates real-world complexity (see Fig. 2(b)). Specifically, DAT provides 24 city-scale scenes, full-fidelity simulations of drone dynamics, and a lightweight tool that can be integrated into any 3D scene to enable human-like target behaviors. It also offers a digital twin tool that can generate unlimited 3D scenes from real-world environments, enabling unlimited scene expansion. Second, we propose a novel drone VAT with reinforcement learning method (called GC-VAT), aiming to improve adaptability in complex and diverse scenarios. Specifically, we design a Goal-Centered Reward to provide precise feedback across viewpoints, enabling the agent to expand perception range through unrestricted perspectives. Besides, we propose qualitative and theoretical methods to analyze the effectiveness of our reward. In addition, inspired by curriculum learning wang2024efficienttrain++ ; scirobot_distill ; 10669618 , we propose a Curriculum-Based Training strategy that progressively improves agent performance in complex environments. Our contributions are summarized as follows:
1) A comprehensive drone active tracking benchmark. We present DAT benchmark, featuring high-fidelity dynamics, 24 city-scale scenes, and tools for simulating human-like target behaviors and unlimited scenes generation, enabling rigorous algorithm validation. 2) A novel drone active tracking method. We propose GC-VAT, which leverages a Goal-Centered Reward function and a Curriculum-Based Training strategy to enhance drone tracking performance in complex and dynamic environments. Besides, we propose qualitative and theoretical methods to analyze the effectiveness of our reward. 3) Extensive experimental validation. Experiments on simulator and real-world images validate DAT usability and GC-VAT effectiveness, with GC-VAT achieving a Tracking Success Rate of approximately 72% on the simulator.
Figure 2: Statistics and simulator component examples of DAT. (a) Statistics on 7 complexity aspects in DAT scenes. (b) Example scenes of DAT. (c) Diverse behaviors of targets. (d) Examples of the tracking targets. More details can be found at https://github.com/SHWplus/DAT_Benchmark.
@@ 2 Task Definition of Drone Active Tracking
DAT task seeks to train a model to control a drone for active target tracking in dynamic environments (see Fig. 1). Using visual and motion sensor data, the model learns actions to keep the target centered in view, ensuring robust performance across diverse scenarios.
Observation spaces. The target is initially positioned at the center of the field of view, and the observation space comprises data acquired from sensors (e.g., RGB images with $84\times 84$ resolution).
Action spaces.
The action space can be either discrete or continuous. A discrete space defines a set of predefined drone maneuvers, whereas a continuous space allows direct control over the velocity.
Success criterion of DAT task. We define a success criterion when the model can keep the target object, which is initially located at the center of view, in the middle of the image for a long duration.
Challenges. Open-world drone active tracking is challenging due to limited data and high risks of trial-and-error in the real world, necessitating complex simulation environments. Additionally, the complexity and dynamics of open-world scenes further demand robust agent performance.
@@ 3 DAT Benchmark with Diverse Settings
We develop DAT, including 24 city-scale scenes built by an unlimited scene generation tool, high-fidelity drone dynamics simulation, and a versatile pipeline for producing human-like target behaviors.
@@ 3.1 Diverse Scene Construction
Digital twin tool. Users can select any region from OpenStreetMap osm to obtain countless scenes using our tool. Specifically, it generates a high-precision road network with traffic lights and rules, and it converts elevation and vegetation data into 3D assets placed in the scene. Moreover, all assets in the generated scene are editable, allowing for data augmentation. See Appendix B for details.
Scene construction. Based on the above tool, we construct 6 outdoor scenes under 4 weather conditions, modeling 7 real-world complexities. Specifically, the scene area, building density, and color richness depict the complexity of the visual background. Road density and terrain density affect the target behaviors. The tree density and tunnel density measure the level of visual occlusion. As shown in Fig. 2(a), six scenarios exhibit unique and realistic complexity across the seven aspects:
@@
•  Village scene spans 1.4 square kilometers. This scene features a mountain density of 20.1 and a tunnel density of 6.3, requiring the agent to predict the target’s movement when it is fully obscured.
•  Downtown scene covers 0.8 square kilometers. It includes complex road elements and high building density of 304.9, challenging the agent’s tracking accuracy and obstacle avoidance abilities.
•  Lake scene encompasses 1.6 square kilometers. The density of road elements is 68.1, and the richness of background colors is 5, challenging the robustness across varying features and colors.
•  Farmland scene covers an area of 0.7 square kilometers. The color richness is 5 and multiple color patches, challenging the agent’s adaptability to multi-color environments.
•  Desert scene covers 1.1 square kilometers. It includes a mountain density of 37.1 and a road density of 31.0. Some roads are covered by sand, testing the agent’s adaptability to such conditions.
Four weather conditions are designed to test the agent’s cross-domain adaptability. Foggy reduces visibility, night reduces brightness, and snow alters the color. The above 24 scenes (see Fig. 2(b)) can fully measure the agent tracking performance. See Appendix B for scene construction details.
@@ 3.2 Various Trackers and Targets Construction
Drone Active Tracking in the real world involves diverse targets depending on tasks. DAT provides diverse targets with human-like behaviors and enables high-fidelity tracker dynamics simulation.
Tracker. DAT benchmark supports two tracker types: drones and ground robots. The drone used is the DJI Matrice 100 m100 , equipped with a 3-axis gimbal, allowing for precise camera adjustments. Unlike simpler kinematic models in dvat and methods that ignore the dynamicse2e , DAT leverages webots webots to simulate the drone’s full dynamics, including mass, inertia, aerodynamics, and the response and jitter of the gimbal, closely matching real drones. See Appendix B for details.
Targets. DAT includes five categories of targets: automobile, motorbike, pedestrian, wheel robot, and legged robot, with a total of 24 tracking targets (see Fig. 2(d)). See Appendix B for details.
Target Management. We propose a novel pipeline to simulate realistic target behavior. Specifically, DAT first utilizes road networks generated by the tools described in Section 3.1, and directly integrates them with the SUMO traffic simulator sumo . Then, random trajectories are assigned to each vehicle, with SUMO managing its motion. To bridge the gap between simulation and visualization, we implement a controller that translates motion data into human-like driving behaviors for 3D vehicles (see Fig. 2(c)). Our controller also adheres to traffic rules and can simulate phenomena such as traffic light waits and traffic jams. Even better, the controller can be applied to any 3D scene.
@@ 4 VAT with Reinforcement Learning
In this paper, we primarily focus on visual active tracking (VAT), a core task within DAT benchmark. We propose a drone visual active tracking with reinforcement learning method called Goal-Centered-VAT (GC-VAT), aiming to improve the performance of tracking targets in complex scenes. As shown in Fig. 1, we model drone active tracking as a Markov Decision Process (MDP) and train a Drone Agent capable of adapting to unrestricted viewpoint conditions to track a target in the open scene.
@@ 4.1 MDP for Drone Active Tracking
Action $\boldsymbol{\mathcal{A}}$ is a set of discrete actions, including forward, backward, leftward, rightward, turn left, turn right, and stop movements. At each time step, the Drone Agent selects an action $a_{t}\in\boldsymbol{\mathcal{A}}$ based on the state $s_{t}$ and actively controls the camera movement.
@@
Reward $\mathcal{R}(s_{t},a_{t})$ is the reward function. The goal-centered rewards are given in Section 4.2.
Network structure of Drone Agent. Similar to previous works e2e ; advat+ , we select a backbone architecture consisting of a CNN followed by a GRU network gru (see Appendix C.2).
Key Challenges in Drone Active Tracking. In open-world environments, drones face unpredictable target behaviors and frequent occlusions. Designing a single reward that encourages diverse and robust tracking actions is extremely difficult. To address this, we propose a goal-centered reward in Section 4.2. Moreover, given the vast observation space, discovering successful policies is non-trivial. To facilitate efficient learning, we introduce a curriculum-based training strategy in Section 4.4.